OpenStudy (anonymous):

find the indicated outputs for f(x) = 4xsecond power - 5x f(0) = f(-1) = f(2) =

OpenStudy (anonymous):

you just plug in the x values that are givin to you ( 0, -1, 2) into the x spaces in the equation. so the first one (f(0) = would be f(0)= 4(0)^2 =5(0) and simplify

OpenStudy (anonymous):

I confused on the last one could you help me

OpenStudy (anonymous):

th f(2) one?

OpenStudy (anonymous):

i got it yes it was the f(2) the answer I got is 6